Here are the dimensions of three rectangles. Which one of them is most square?
114’ × 99’ 455’ × 494’ 284’ × 265’

Found 2 solutions by shree840, Alan3354:
Answer by shree840(260)

99-114)/114=15/114=0.13
494-455)/494=.07
284-265)/284=.066
third one is most square ans

Answer by Alan3354(69443)

Rectangles are either squares or not squares.
There is no "most square."
The other tutor showed which one is most nearly square.